1. A balloon at an altitude of 30 km is filled with helium gas at a pressure of 10-2 atm and a temperature of 250 K. Estimate roughly the number of helium atoms hitting each square nanometer of the inner surface of the balloon per nanosecond.
Answer #1

From Kinetic theory of gases

vrms=(3RT/M)^.5

=1248.549m/s

P=dRT/M

d=density

10^-2*10^5=d*8.314*250/4*10^-3

d=1.9*10^-3kg/m^3

for finding number of molecules/s volume of balloon will be required